OpenStudy (anonymous):

cos(a+B) - cos(a-B).... i'm getting 0.... i feel like it's wrong

geerky42 (geerky42):

Remember, \(\cos(\alpha \pm \beta) = \cos\alpha\cos\beta \mp \sin\alpha\sin\beta\)

OpenStudy (anonymous):

=cosa*cosb - sena*senb - (cosa*cosb + sena*senb) = -2*sena*senb